OpenStudy (heyitslizzy13):

How do I solve 2x^2-3x-5=0 by factoring?

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/2x~2-3x-5=0/ this might help you

OpenStudy (hayleymeyer):

no comment.

OpenStudy (anonymous):

no comment needed ma'am

OpenStudy (hayleymeyer):

XDD how old are you?

OpenStudy (anonymous):

im 18 lol

OpenStudy (hayleymeyer):

Dammmmmmmmmmmmmmmmmmn!

OpenStudy (heyitslizzy13):

i'm kind of more confused :/ is there any way you can explain it?

OpenStudy (anonymous):

im in pre calc lol

OpenStudy (anonymous):

yes give me a second ill give you my notes

OpenStudy (anonymous):

first read this hayley https://answers.yahoo.com/question/index?qid=20080707161908AA6NmWp

OpenStudy (hayleymeyer):

i am 16, i wouldve thought older guys would be mean to us young girls xD

OpenStudy (heyitslizzy13):

thank you

OpenStudy (anonymous):

naw lol im from the south i respect girls women

OpenStudy (anonymous):

yw may i have a medal plz

OpenStudy (hayleymeyer):

we need more people like you in St. Louis <33

OpenStudy (anonymous):

lol you find them there rarley but just keep searchin lol follow your heart not what evrybody else says love is better then words

OpenStudy (anonymous):

@Hayleymeyer

OpenStudy (hayleymeyer):

* blows kisses* XD. you've made my day.

OpenStudy (anonymous):

bows thank you thank you very much lol

OpenStudy (hayleymeyer):

:D you should join the mathmstics chat

OpenStudy (hayleymeyer):

Mathmatics*

OpenStudy (heyitslizzy13):

i don't understand

OpenStudy (hayleymeyer):

@Tazmaniadevil

OpenStudy (hayleymeyer):

i called him bacckkk. :3.

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

(2x - 5)(x + 1) = 0 2x - 5 = 0 or x + 1 = 0 2x = 5 or x = -1 x = 5/2 or -1 ok what dont you understand

OpenStudy (heyitslizzy13):

where is the squared?